# Follow Everything: A Leader-Following and Obstacle Avoidance Framework with Goal-Aware Adaptation

### v1

@@ Abstract
Robust and flexible leader-following is a critical capability for robots to integrate into human society. While existing methods struggle to generalize to leaders of arbitrary form and often fail when the leader temporarily leaves the robot’s field of view, this work introduces a unified framework addressing both challenges. First, traditional detection models are replaced with a segmentation model, allowing the leader to be anything. To enhance recognition robustness, a distance frame buffer is implemented that stores leader embeddings at multiple distances, accounting for the unique characteristics of leader-following tasks. Second, a goal-aware adaptation mechanism is designed to govern robot planning states based on the leader’s visibility and motion, complemented by a graph-based planner that generates candidate trajectories for each state, ensuring efficient following with obstacle avoidance. Simulations and real-world experiments with a legged robot follower and various leaders (human, ground robot, UAV, legged robot, stop sign) in both indoor and outdoor environments show competitive improvements in follow success rate, reduced visual loss duration, lower collision rate, and decreased leader-follower distance. Visit the website for the video and code.
@@ I Introduction
Recent advances in visual perception, robot navigation, and large models have enabled robots to follow leaders with increased safety and intelligence [1, 2, 3]. Serving as powerful assistants, robots can follow explorers to carry more supplies, guide customers and introduce products in shopping malls, help police track and capture suspects, and more [4, 5, 6].
While existing research on leader-following typically focuses on following a specific pedestrian detected using lightweight camera-based [7], LiDAR-based [8], or multimodal-based [9, 10] approaches, often assisted by Kalman filters [11], we argue that a robot’s following capability should extend beyond this limited scope, enabling it to follow leaders of any shape or type, including humans, animals, vehicles, irregular objects, and even other robots.
@@
To support this, instead of relying on conventional detection models designed for specific categories, this work adopts a segmentation model, SAM2 [12], which enables flexible selection of leader masks with arbitrary shapes. This shift in logic enables anything to be a leader, regardless of its shape or type. Unlike the work [13], which only uses SAM2 with RGB image input, our method better accommodates the leader-following task, where the depth image is accessible and the distance between the leader and the robot fluctuates within a certain range. A distance frame buffer is introduced (see Sec.II) on top of the existing model [12], enhancing tracking robustness, particularly when the leader re-enters the robot’s field of view (FOV) after a temporary loss.
However, both detection- and segmentation-based methods are inherently limited by the FOV constraints of onboard sensors [14]. As a result, once the leader moves beyond the sensor’s FOV, the robot may fail to detect the leader, leading to awkward hesitation or inaction [15]. To address this limitation, some studies [16] have incorporated sonar sensors to detect the leader’s footsteps; however, this approach tends to perform poorly in crowded environments. Other methods rely on the leader carrying lightweight beacons, such as infrared antennas [17] or wave emitters [18, 19], but requiring the leader to bear additional equipment is often impractical.
While these methods do help alleviate the issue of lossing the leader by improving the perception robustness, some instances are inevitable [20]. For example, when a mobile robot follows a human walking out of a room, the human will naturally disappear from the robot’s current FOV after exiting. Similarly, when a ground robot follows a UAV, the UAV may fly over obstacles that the ground robot must detour around, making it easy to lose visual contact during the process [13]. Therefore, we argue that it is impractical to rely solely on the perception module to handle such cases. The planning module should work in conjunction with perception and be capable of actively re-locating the leader after temporary loss, thereby enhancing the robustness of the following behavior— for instance, by adaptively adjusting the robot’s speed when the leader moves quickly or becomes distant, and maintaining an appropriate safety distance to reduce the risk of losing the leader. In addition, when the leader approaches the robot, the robot should be able to retreat accordingly to avoid getting too close. This helps prevent both failure in capturing the leader’s mask due to excessive proximity and the collision risk when the leader steps backward without being aware of the robot’s position.
@@
Figure 2: Illustration of the proposed leader-following and obstacle avoidance framework. Given a leader prompt and an RGB image, SAM segments the leader mask with the aid of both a memory buffer and a distance frame buffer. The segmented leader is then filtered from raw laser point clouds to produce a leader point set, which is passed to the goal-aware adaptation module to determine the current goal state and relevant parameters. Meanwhile, a costmap is generated and sent to the planning module, which computes the optimal trajectory based on the selected goal state and parameters.
While most existing approaches overlook the issues mentioned above, simply treating the leader’s last known position as a waypoint and planning motions using PID-based controllers [13], MPC-based controllers [21], sampling-based trajectories [22], optimization-based trajectories [23], or learning-based approaches [24], the goal-aware adaptation proposed in this work systematically considers these factors within an integrated framework (see Sec.III).
Five states are defined within the goal-aware adaptation framework to address various leader-robot interaction scenarios: (a) The following state is activated when the leader is within the robot’s field of view, and the robot’s goal is a set of points on a variable-radius circle centered on the leader. The environment is modeled as a topological graph to generate and optimize candidate trajectories, whose endpoints can slide along the circular goal set. The robot’s maximum speed is determined by its relative state to the leader, while the safety distance (circle radius) is dynamically adjusted based on the leader’s motion uncertainty. (b) The chasing is activated when the leader is visible but at a considerable distance, and the robot’s goal is a set of lines pointing toward the leader. The robot then moves at its maximum allowable speed to minimize the risk of losing the leader. (c) The retreating state is activated when the leader moves toward the robot. The goal remains on the circular point set, but additional constraints ensure that the robot maintains its orientation toward the leader while retreating to a safe distance. (d) The planning state is triggered when the leader leaves the robot’s FOV. The robot uses the leader’s last known state as a temporary goal and computes a retracing trajectory via the topological graph. (e) The switching state allows dynamic leader replacement. External prompts, interpreted by a large language model [12], generate a new leader mask, enabling the robot to follow a newly assigned leader.
@@
In summary, this work presents a unified framework for robustly leader-following and obstacle avoidance. The main contributions are as follows:
@@
  • A segmentation module with a distance frame buffer is introduced to enable anything to be a leader.
@@
II Leader Segmentation with Distance Frame Buffer
Given the initial RGB image $I_{0}$ and a set of leader point prompts (often manually selected on the image; see Fig.3a), the segmentation model, EVF-SAM [12], automatically expands the prompts to segment the full leader, producing an embedding $\eta_{0}$ and its corresponding mask $\phi_{0}$. At each subsequent timestep $t$, the top $n$ historical leader embeddings with the highest confidence scores are updated and maintained in a temporal memory buffer:
$\mathcal{B}^{T}=\{\eta_{i},\eta_{j},\eta_{k},...\}_{n}$ (1)
These embeddings, along with the current image $I_{t}$, are utilized to generate the current leader embedding and mask:
$\eta_{t},\phi_{t}=\text{SAM}(I_{t},\mathcal{B}^{T})$ (2)
This temporal memory buffer performs well when the leader remains continuously within the robot’s FOV. However, as illustrated in Fig.1(a), the leader may be temporarily lost when the robot detours around obstacles. Just before this loss, the leader often appears at the edge of the robot’s FOV, resulting in partial and low-quality embeddings. Despite their poor quality, these embeddings may still dominate the memory buffer due to their relatively higher confidence scores, since embeddings that capture the leader’s full body shape differ significantly from the current appearance of the leader in the image [13]. This may mislead the robot to bootstrap toward incorrect directions. As a result, when the leader reappears later, the robot may fail to retrace the leader accurately, even if the complete leader is in front of it.
@@ II Leader Segmentation with Distance Frame Buffer
Figure 3: Illustration of goal-aware adaptation and graph-based planning in a demonstration. (a) SAM segments the leader after manually selecting a few leader points. Chasing is activated as the leader is far away. (b) Following is triggered when the robot gets closer, maintaining a safe distance. (c) Planning is activated when the leader exits the robot’s FOV, guiding the robot to the last known leader pose. (d) Following resumes once the leader is visible again. (e) Retreating is triggered when the leader steps back. (f) A new prompt activates Switching, followed by Chasing as the new leader is distant.
To address this issue, the distance frame buffer $\mathcal{B}^{D}$ is introduced. Unlike the temporal memory buffer, which operates along the time dimension, the distance frame buffer leverages spatial information enabled by depth image input. In leader-following tasks, the distance $d_{i}$ between the robot and the leader varies dynamically due to speed differences or detours. For each relative distance segment (with a margin $\Delta d$), the leader embedding with the highest confidence score is stored and updated in the distance frame buffer:
$\begin{gathered}\mathcal{B}^{D}=\{\eta_{i},\eta_{j},\eta_{k},...\}_{n}\\ \eta_{t},\phi_{t}=\text{SAM}(I_{t_{i}},\mathcal{B}^{T},\mathcal{B}^{D})\end{gathered}$ (3)
By incorporating both temporal and spatial memory, this mechanism significantly improves SAM’s ability to retrace and re-identify the leader after temporary visual loss.
Aligning the leader mask $\phi_{t}$ with the corresponding depth image, the leader can be extracted as a point set $\mathcal{P}_{leader}=\{\mathbf{p}_{1},\mathbf{p}_{2},...,\mathbf{p}_{m}\}$, where each point $\mathbf{p}_{i}=(x,y,z)$. These points are then used to filter the point cloud transformed from the laser scan, and the remaining points form a costmap $\mathcal{M}_{t}$.
III Goal-aware Adaptation and Graph-based Trajectory Planning
III-A Parameters Adaptation
During planning, two key parameters related to the leader’s state influence the robot’s behavior: the maximum speed $V_{t}^{\text{max}}$ and the safe distance $D_{t}$.
The robot’s maximum speed is determined by the relative distance between the robot and the leader, as well as the leader’s estimated velocity. The leader’s position $\bar{p}_{l}$ is obtained by averaging the segmented leader point set $\mathcal{P}_{leader}$, and the leader’s velocity $\bar{v}_{l}$ is estimated based on the positional difference over a fixed time interval. The maximum robot speed during the current planning is then computed as:
$V_{t}^{max}=\text{Clip}\left(\alpha_{1}|\bar{v}_{l}|+\alpha_{2}|\bar{p}_{l}-p_{r}|,\,0,\,V_{max}\right)$ (4)
where $\alpha_{1}$ and $\alpha_{2}$ are predefined scaling factors, $p_{r}$ is the robot’s position, ||$\cdot$|| ⋅ | denotes the vector norm, and $V^{max}$ is the maximum velocity the robot can physically execute. The $\text{Clip}(\cdot)$ function ensures the computed value respects the robot’s dynamic constraints.
To ensure safe interaction, the robot maintains a larger safe distance $D_{t}$ when the leader’s motion is more uncertain. A Kalman filter is used to estimate the leader’s state, where the Normalized Innovation Squared (NIS) [11] value serves as an indicator of prediction uncertainty. Based on this, the safe distance is dynamically adjusted as:
$D_{t}=\text{Clip}\left(\alpha\text{NIS},\,D^{min},\,D^{max}\right)$ (5)
where $D^{min}$ and $D^{max}$ are predefined lower and upper bounds, and $\alpha$ is a scaling factor.
III-B Goal States Adaptation and Trajectory Planning
The transition between states is governed by conditions including whether the leader is within the robot’s FOV, whether it lies within the costmap, the distance between the robot and the leader, and whether a new leader is assigned.
The chasing state (see Fig.3a) is activated when the leader is within the robot’s FOV but outside the current costmap range. In this state, the robot’s primary task is to navigate around nearby obstacles and chase the leader as quickly as possible. Following prior work [25, 26], adjacent obstacle points in the costmap are clustered into obstacle groups, each outlined by a boundary. Each obstacle group with its boundary forms a node $\mathcal{N}_{i}$ in the topological graph. For any two boundaries with indices $i$ and $j$, their shortest collision-free connection is defined as an edge $\mathcal{E}_{i,j}$.

### v3

# Follow Everything: Goal-Aware Adaptation and Graph-Based Planning Towards Arbitrary Leader Following

@@ Abstract
Enabling robots to robustly follow leaders supports tasks such as carrying supplies or guiding customers. While existing methods often fail to generalize to arbitrary leaders, such as humans or other robots, and struggle when the leader temporarily leaves the robot’s field of view, this work presents a unified framework to address both challenges. First, a segmentation model replaces traditional category-specific detection models, allowing the leader to be of any shape or type. To improve robustness, a distance frame buffer is designed to store high-confidence leader embeddings across distance intervals, accounting for the unique characteristics of leader-following tasks. Second, a goal-aware adaptation mechanism is designed to govern robot planning states based on the leader’s visibility and motion, complemented by a graph-based planner that generates candidate trajectories for each state, ensuring efficient following with obstacle avoidance. Simulations and real-world experiments with a legged robot follower and diverse leaders (e.g., humans, UAVs, legged robots, and stop signs) in indoor and outdoor settings demonstrate an improved follow success rate of 75.1%, a reduced visual loss of 13.1%, a fewer collisions of 65.1%, and a shorter average distance of 0.4 m. Visit the website for the video and code.
@@ I Introduction
Recent advances in visual perception, robot navigation, and large models have enabled robots to follow leaders with increased safety and intelligence [1, 2, 3]. Serving as powerful assistants, robots can follow explorers to carry more supplies, guide customers and introduce products in shopping centers, help police patrolling, etc. [4, 5, 6]. Although numerous works have been proposed to improve leader following stability [7, 8, 9] or robot movement smoothness [10, 11, 12, 13, 14, 15], all overlook a unique challenge in the leader-following task: the leader may temporarily leave the robot’s field of view (FOV), as shown in Fig. 1(a), which causes existing methods to fail in finding a trajectory without knowing the leader’s position or in not recognizing the leaders when they reappear. This paper breaks down this challenge from both the perception and planning perspectives and proposes the distance frame buffer, goal-aware adaptation, and graph-based planner to address it.
@@
The challenge on the perception side concerns whether the leader can be identified as the same leader when they leave and then reappear in the follower’s FOV. One specific approach for leader-following tasks is to have the leader carry a beacon, such as infrared antennas [16] or wave emitters [17, 18], but requiring the leader to carry additional equipment is often impractical and not generalizable. Therefore, another widely used solution is to leverage detection models based on cameras [7], LiDAR [8], multimodal sensors [9], sonar [19], and so on. However, since their training labels correspond to object categories rather than object features, when the leader (typically a pedestrian) remains in the follower’s FOV, its ID remains stable due to the help of the Kalman Filter [20]. However, when the leader leaves and reappears, a new ID will be assigned to it. This leads to the expected leader having the original ID permanently lost.
@@
Figure 2: Illustration of the proposed follow everything framework. Given a leader prompt and an RGB image, the leader mask is segmented with the aid of both a memory buffer and a distance frame buffer. The segmented leader is then filtered from raw laser point clouds to produce a leader point set, which is passed to the goal-aware adaptation to dynamically adjust parameters and provide goal sets and constraints according to the follower-leader interaction. Given the costmap, goal sets, and constraints, the graph-based planner searches, optimizes, and selects an optimal trajectory for the follower to execute.
An alternative solution is segmentation models [21], which use features as labels. After the initial feature is provided, the model continuously tracks and bootstraps-updates the leader feature, allowing the leader to be reidentified when it reappears [10]. Not limited to categories, this type of model also enables anything with an arbitrary shape to be the leader, such as humans, vehicles, irregular objects, or even other robots. However, a limitation of existing segmentation datasets is that they often involve a fixed camera position (follower) with the leader’s movement, location and size constantly changing in the video, where bootstrapping is beneficial. However, in the leader-follower task, the follower is also in motion, making it easy for the leader to leave the follower’s FOV, as illustrated in Fig. 1(a). Just before leaving, the leader often appears on the edge of the robot’s FOV, resulting in partial and low-quality features. Despite their poor quality, these features can still dominate the bootstrapping process, since the older features that capture the full-body shape of the leader differ significantly from the current appearance of the leader [10]. This can mislead the robot to bootstrap in the wrong direction. As a result, when the leader reappears later, the robot may not be able to retrace the leader accurately, even if the full leader is in front of it. To address this issue, this work introduces the distance frame buffer to SAM2 (see Sec. II-B), which divides the distance ranges between the follower and leader into equal intervals and maintains the historical leader feature with the highest confidence in each distance interval, greatly improving the ability to retrace the leader after it reappears.
The challenge on the planning side concerns obstacle avoidance and robust motion planning when the relationship between the leader and the follower is dynamic. While most existing approaches simply treat the leader’s position as a target and use PID [10], MPC [11], sampled trajectories [12], optimized trajectories [13], or learning-based methods [15] to plan trajectories for the follower (robot), they often lack robustness and fail to consider the various interactive states that arise from the leader-follower relationship. For example, (a) when the leader is within the follower’s FOV and close by, the main goal of the follower is to avoid obstacles while following, a scenario that frequently occurs when a legged robot follows a UAV leader, as shown in Fig. 6, where the UAV can fly over short obstacles that the ground robot must detour around. (b) When the leader is within the follower’s FOV but at a greater distance, the main goal of the follower is to quickly approach the leader, which requires the robot to maintain high speed and choose optimal time trajectories to avoid increasing the risk of losing the leader due to excessive distance, as shown in Fig. 8. (c) When the leader retreats, the follower should also retreat accordingly while maintaining a safe margin to avoid collisions with the leader or losing sight of it due to too-close proximity, as shown in Fig. 5d. (d) When the leader leaves the follower’s FOV, the follower should intentionally approach the last known position and orientation of the leader to recover it as quickly as possible, as illustrated in Fig. 9f-g. (e) A robust model should also be able to replace the leader if necessary, improving the practicality. To realize these robust following capabilities, this work designs a goal-aware adaptation (see Sec.II-D) under the proposed follow everything framework, which dynamically adjusts the robot’s speed and the desired safe distance between the leader and the follower, and adaptively determines the follower’s state to provide unique goals and constraints for the subsequent trajectory planner.
As for the planner, the primary concern is to find a time-optimal trajectory for the follower while adapting to the goals and constraints provided by the goal-aware adaptation. Building on previous work [22, 23], the environment can be modeled as a graph, with obstacles clustered as nodes and their shortest connections as edges. Given that each node has different detour directions, a series of candidate trajectories with different meanings [24, 25] can be generated. Besides kinematic constraints, the goal sets and unique constraints provided by goal-aware adaptation are incorporated into the optimization. The trajectory with the shortest time to reach the goal is selected as the best trajectory to execute. This graph-based planner (see Sec. II-C) enables the follower to approach the leader efficiently, rather than simply replicating their historical paths, as shown in Fig. 1(b).
@@
In summary, this work presents a unified framework for robust leader-following and obstacle avoidance. The main contributions are as follows:
@@
  • A segmentation module with a distance frame buffer is introduced to enable everything to be a leader.
@@
Figure 3: Illustration of goal-aware adaptation and graph-based planner in a demo. (a) The leader is segmented after selecting points. Chasing is activated as the leader is far away. (b) Following is triggered when the robot gets closer, maintaining a safe distance. (c) Planning is activated when the leader exits the robot’s FOV, guiding the robot to the last known leader pose. (d) Following resumes once the leader is visible again. (e) Retreating is triggered when the leader steps back. (f) A new prompt "follow the person on the left side" activates Switching, followed by Chasing as the new leader is distant.
II Methodology
II-A Problem Formulation
Given the RGB image $I_{t}$ at timestamp $t$, the leader is segmented into an embedding $\eta_{t}$ and a mask $\phi_{t}$ with the assistance of a temporal memory buffer $\mathcal{B}_{t}^{T}$ and a distance frame buffer $\mathcal{B}_{t}^{D}$. Using the depth image, the leader mask $\phi_{t}$ is transformed into a point set $\mathcal{P}^{leader}_{t}=\{\textbf{p}_{1},\textbf{p}_{2},...,\textbf{p}_{m}\}$, where each point $\textbf{p}_{i}=(x,y,z)$ represents the 3D coordinates of the leader. The leader’s position $\bar{\textbf{p}}^{l}_{t}$ is obtained by averaging the point set $\mathcal{P}^{leader}_{t}$, and the leader’s velocity $\bar{\textbf{v}}^{l}_{t}$ is estimated based on the positional difference over a fixed time interval. Given the robot’s position $\textbf{p}^{r}_{t}$ and the leader’s states $\bar{\textbf{p}}^{l}_{t}$ and $\bar{\textbf{v}}^{l}_{t}$, the goal-aware adaptation mechanism determines the robot’s maximum speed $V_{t}^{max}$ and the safe distance $D_{t}$ between the robot and the leader. Additionally, the robot’s goal is provided as a series of sets $\{\mathcal{G}_{t}^{0},\mathcal{G}_{t}^{1},...,\mathcal{G}_{t}^{n}\}$, each associated with the desired orientations. A topological graph is constructed based on a costmap $\mathcal{M}_{t}$, where obstacle groups are defined as nodes $\mathcal{N}_{i}$ and the shortest connections between groups are defined as edges $\mathcal{E}_{i,j}$. This graph generates several generalized trajectories $\mathcal{T}_{i}$, which are then transformed into normal trajectories $\tau_{i}$. By optimizing these trajectories with constraints imposed by the robot’s kinematics, goal sets $\mathcal{G}_{t}^{i}$, maximum speed $V_{t}^{max}$, and safe distance $D_{t}$, the optimal trajectory is selected and executed.
II-B Leader Segmentation with Distance Frame Buffer
Given an initial RGB image $I_{0}$ and a set of leader point prompts (often manually selected; see Fig. 3a), the segmentation model [21] expands the prompts to the entire leader, producing a leader embedding $\eta_{0}$ and a leader mask $\phi_{0}$. Then, at each subsequent timestep $t$, the top $n$ historical leader embeddings with the highest confidence scores are maintained and updated in a temporal memory buffer $\mathcal{B}_{t}^{T}$:
$\mathcal{B}^{T}_{t}=\{\eta^{1},...\}_{n_{1}},\;\eta^{i}=\mathop{\arg\max}\limits_{\forall\eta_{t}\;\cap\;\eta_{t}\notin[\eta^{1},...,\eta^{i-1}]}\;S(\eta_{t}),$ (1)
where $\forall\eta_{t}$ contains all historical embeddings and $S(\eta_{t})$ is the confidence score. The distance frame buffer $\mathcal{B}_{t}^{D}$ considers the spatial dimension. It has a length of $n_{2}$, with each element $\eta^{i}$ storing the leader embedding with the highest confidence score from historical leader embeddings within the distance range $[(i-1)\Delta d,i\Delta d]$ between the follower and the leader:
$\mathcal{B}_{t}^{D}=\{\eta^{1},...\}_{n_{2}},\eta^{i}=\mathop{\arg\max}\limits_{\forall\eta_{t}\;\cap\;d(\eta_{t})\in[(i-1)\Delta d,i\Delta d]}\;S(\eta_{t}),$ (2)
where $\Delta d$ is a distance interval, and the distance between the follower and the leader is defined as $d(\eta_{t})=|\bar{\mathbf{p}}_{t}^{l}-\mathbf{p}_{t}^{f}|$. Given the current image $I_{t}$ and the two buffers, the current leader embedding $\eta_{t}$ and mask $\phi_{t}$ can be generated:
$\eta_{t},\phi_{t}=\text{Segmentation}(I_{t},\mathcal{B}^{T}_{t},\mathcal{B}^{D}_{t}).$ (3)
The introduction of the distance frame buffer enables the segmentation model to isolate embeddings based on distance, reducing the contamination of the entire buffer when bootstrap occurs in the wrong direction, such as when the leader is partially observable or in the process of leaving the follower’s FOV. Its effectiveness can be seen in Table. LABEL:table, where our Follow Everything achieves the highest follow success rate metric compared to Alaa and FE-N-DFB.
II-C Graph-based Planner
The graph-based planner is based on time-optimal optimization [26] and provides a unified framework to incorporate various constraints of goal-aware adaptation, making the decision-making of the robot consistent even when motion states frequently alternate. Filtering the leader points $\mathcal{P}_{leader}$ from the laser scan, the remaining points form a binary costmap $\mathcal{M}_{t}$. Following prior work [22, 23], adjacent obstacle points in $\mathcal{M}_{t}$ are clustered into obstacle groups, each outlined by a boundary. A topological graph is established with the obstacle groups as nodes $\mathcal{N}_{i}$. For any two boundaries with indices $i$ and $j$, their shortest collision-free connection is defined as an edge $\mathcal{E}_{i,j}$ (see Fig. 3a) of the graph. Given the goal point or goal sets $\{\mathcal{G}_{t}^{0},\mathcal{G}_{t}^{1},...\}$ from the goal-aware adaptation, which will be detailed later, several generalized trajectories can be generated over the topological graph using depth-first search, with each one $\mathcal{T}$ following the form:
[…]
Since each obstacle node $\mathcal{N}_{i}$ can be detoured in clockwise or counterclockwise directions, a generalized trajectory $\mathcal{T}$ containing $n$ nodes can produce up to $2^{n}$ distinct trajectories $\tau$. To eliminate redundant trajectories that represent the same detour behavior, the trajectory homotopy classes [24, 25] are applied. The remaining trajectories are then optimized in parallel. For each trajectory $\tau=\{\textbf{p}_{0},\textbf{p}_{1},...,\textbf{p}_{m}\}$, it follows the time-optimal optimization framework:
$\tau$ $=\mathop{\arg\min}\limits_{\{\textbf{p}_{1},...,\textbf{p}_{m}\}}\sum_{t=1}^{m}||\textbf{p}_{i}-\textbf{p}_{i-1}||_{2}/V_{t}^{max}.$ (5)
[…]
$\quad\quad\mathcal{H}(\textbf{p}_{i},\textbf{p}_{i+1},\textbf{p}_{i+2})=0,\quad\mathcal{A}(\textbf{p}_{i},\textbf{p}_{i+1},\textbf{p}_{i+2})\leq 0,$ (7)
$\quad\quad\mathcal{I}(\textbf{p}_{m},\mathcal{G}_{i})\leq 0,\quad\mathcal{V}(\textbf{p}_{i},\textbf{p}_{i+1},V_{t}^{max})\leq 0,$ (8)
where $\mathcal{O}_{dynamic}(\cdot)$ and $\mathcal{O}_{static}(\cdot)$ represent the constraints for avoiding dynamic [22] and static [23] obstacles, respectively. $\mathcal{H}(\cdot)$ and $\mathcal{A}(\cdot)$ denote the kinematic and acceleration constraints [26], respectively. $\mathcal{I}(\cdot)$ and $\mathcal{V}(\cdot)$ are the goal and velocity constraints, both of which are provided by the goal-aware adaptation and will be detailed in the next subsection.
All the trajectories are evaluated, and the one $\tau^{*}_{t}$ with the lowest cost is selected and executed by the robot:
$\begin{gathered}\tau^{*}_{t}=\mathop{\arg\min}\limits_{\tau}f(\tau^{*}_{t-1},\tau)g(\tau),\\ f(\tau^{*}_{t-1},\tau)=(1-\alpha)+\alpha\sum_{i=1}^{l}\frac{val_{i}(\tau^{*}_{t-1})-val_{i}(\tau)}{l},\end{gathered}$ (9)
where $f(\tau^{*}_{t-1},\tau)$ evaluates the similarity between the previous optimal trajectory $\tau^{*}_{t-1}$ and the current candidate $\tau$, and $g(\tau)$ evaluates the time cost of the trajectory, as defined in Equ.(5). In a scenario with $l$ obstacle groups, a trajectory is assigned a homology signature $\{val_{1},val_{2},...,val_{l}\}$, where each $val_{i}\in\{1,-1\}$ indicates whether the trajectory detours the $i$-th obstacle group in a clockwise or counter-clockwise direction, respectively. $\alpha$ is a scaling factor. This design enhances the robustness of trajectory selection, even when the goal sets dynamically change due to goal-aware adaptation.
II-D Goal-aware Adaptation
The goal-aware adaptation governs the goal constraint $\mathcal{I}(\cdot)$ and velocity constraint $\mathcal{V}(\cdot)$ during trajectory optimization, balancing safety in avoiding obstacles and efficiency in following the leader. The transition between goal states is determined by conditions, including whether the leader is within the robot’s FOV, whether it lies within the costmap, the robot-leader distance, and whether a new leader is assigned. See Fig. 2 for possible transitions between states.